# Discriminating image representations with principal distortions

### v1

@@ D.2 Supplemental Figures
Figure 6: Principal distortions for ResNet50 and AlexNet for six example base images (${\mathbf{s}}$, middle row). Each row shows base images with differently scaled additive perturbations of ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{2}$. Distortions $\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}$ are shown in isolation in the top/bottom rows, respectively. The log ratio plot for each image is at the top of the column.
@@
Figure 7: Principal distortions for ResNet50 and AlexNet for base images that have been constructed to have either (A) texture only in the periphery and a blank center (B) texture only in the center and a blank surround and (C) a full image of texture. The log ratio plot for each image is at the top of the column. Panels A and B highlight that the AlexNet architecture is more sensitive to the perturbation with power around the “stuff” of the image (i.e., the non-blank areas), while ResNet50 is more sensitive to a distortion focused on blank areas of the image, and these perturbations are not sensitive to the choice of low contrast (“black”) or high luminance (“white”) for the blank part of the space. Panel C highlights that in the case where there is not obvious blank area of the image, the perturbations become harder to interpret, and this reflects the locally adaptive property of the FI to the input image.
@@
Figure 8: Principal distortions for standard ImageNet trained and Shape Image Net Trained (SIN) AlexNet and ResNet50 architectures for example base images.
@@
Figure 9: Principal distortions for standard ImageNet trained and adversarially trained AlexNet and ResNet50 architectures for example base images.
@@
Figure 10: Example principal distortions for ResNet50 vs. AlexNet comparison where the pixelwise min/max value for the perturbation has not been constrained to avoid clipping once the perturbation is scaled. Similar to the results in the main text, we see that the principal distortion that AlexNet is more sensitive to is focused on the part of the image with high-contrast texture features, while ResNet50 is more sensitive to the perturbation targeting relatively blank parts of the image.

### v2

@@ D.3 Supplemental Figures
Figure SI.3: Principal distortions for ResNet50 and AlexNet for six example base images (${\mathbf{s}}$, middle row). Each row shows base images with differently scaled additive perturbations of ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{2}$. Distortions $\{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\}$ are shown in isolation in the top/bottom rows, respectively. The log ratio plot for each image is at the top of the column.
Figure SI.4: Adding a principal distortion to the base image changes the prediction of one DNN but not the other. Principal distortions were generated for intermediate layers of the AlexNet and ResNet architectures and were computed for a set of 72 base images from the ImageNet validation set (chosen from images included in ImageNet-9, Xiao et al., 2021). For each image ${\mathbf{s}}$ and principal distortion ${\bm{\epsilon}}_{i}$, the distorted images ${\mathbf{s}}+k_{i}{\bm{\epsilon}}_{i}$, with $k_{i}$ varying between 0 and 1, were classified by the DNNs used for the principal distotion generation (AlexNet or ResNet50). The left plot shows the proportion of images for which each DNN had the same classification for the distorted image ${\mathbf{s}}+k_{1}{\bm{\epsilon}}_{1}$ as the base image ${\mathbf{s}}$, for $k_{1}$ varying between 0 and 1. The right plot is the same, but with ${\bm{\epsilon}}_{2}$ and $k_{2}$ in place of ${\bm{\epsilon}}_{1}$ and $k_{1}$.
Figure SI.5: Principal distortions generated for multiple random initializations of AlexNet and ResNet50 consistently separate models by architecture type. A) Log sensitivity ratios of principal distortions for networks trained on standard ImageNet when comparing image representations at multiple layers of AlexNet and ResNet50, using three random initializations of each architecture. Principal distortions were computed for a set of 72 base images from the ImageNet validation set (chosen from images included in ImageNet-9, Xiao et al., 2021), and the mean ratio across these 72 images is plotted (error bars are standard deviation). B) Example base image, principal distortions, and distorted images. The principal distortions are qualitatively similar to those observed from a single architecture in Fig. 3. C) Proportion of images with a changed prediction class (compared to image ${\mathbf{s}}$) for the distorted images ${\mathbf{s}}+k_{i}{\bm{\epsilon}}_{i}$. Results are similar to trends for single seeds of models from Supp. Fig. SI.4.
@@
Figure SI.6: Principal distortions for ResNet50 and AlexNet for base images that have been constructed to have either (A) texture only in the periphery and a blank center (B) texture only in the center and a blank surround and (C) a full image of texture. The log ratio plot for each image is at the top of the column. Panels A and B highlight that the AlexNet architecture is more sensitive to the perturbation with power around the “stuff” of the image (i.e., the non-blank areas), while ResNet50 is more sensitive to a distortion focused on blank areas of the image, and these perturbations are not sensitive to the choice of low contrast (“black”) or high luminance (“white”) for the blank part of the space. Panel C highlights that in the case where there is not obvious blank area of the image, the perturbations become harder to interpret, and this reflects the locally adaptive property of the FI to the input image.
@@
Figure SI.7: Principal distortions for standard ImageNet trained and Shape Image Net Trained (SIN) AlexNet and ResNet50 architectures for example base images.
Figure SI.8: Control experiment dissociating foreground and background from high- and low-spatial frequency regions. (A) Example natural base image, image with a blurred background, and image with a blurred foreground. Images were chosen from ImageNet-9 (Xiao et al., 2021), which contains masks of the “foreground” parts of the image. Images are $224\times 224$ pixels. To construct the blurred variants, a Gaussian filter with $\sigma=5$ is was applied to the full image, and we used the provided image masks to construct the Blur Background (Blur BG) condition by adding the original image foreground to the blurred image background, and the Blur Foreground (Blur FG) condition by adding the original image background to the blurred image foreground. The associated optimal distortions are shown for each image, generated from the layers from AlexNet, ResNet50, SIN-Trained AlexNet, and SIN-Trained ResNet50. In the example image, the distortion AlexNet is most sensitive to (${\bm{\epsilon}}_{1}$) is biased towards higher frequency (non-blurred) parts of the image, while the distortion ResNet50 is most sensitive to (${\bm{\epsilon}}_{2}$) is biased towards low-frequency (blurred) parts of the image. We quantified this observation in (B) by masking the perturbation with the foreground mask, and measuring the average squared value of each perturbation in this region. If a perturbation is localized to the foreground, it will have a high perturbation magnitude on this plot. Each point on the plot corresponds to one of 72 randomly selected images from ImageNet-9 (Xiao et al., 2021), where 8 images were selected from each category, and the box and whisker plots are defined with the median given by the line across the box, the box extends from Q1 to Q3, and the whiskers extend to 1.5 IQR. Although there is a slight bias for distortion ${\bm{\epsilon}}_{1}$ (defined as the perturbation AlexNet is more sensitive to), to have more power in the foreground of the image, this is likely due to the general biases of natural images where often times the background is blurred. If we blur the background, we see that this difference is exaggerated, while if we blur the foreground, the AlexNet perturbation shifts to be less concentrated in the foreground while the perturbation ResNet50 is more sensitive to (${\bm{\epsilon}}_{2}$) becomes more concentrated in the foreground. Points for individual categories are shown in (C), which show the same trends, and results look similar when just including AlexNet and ResNet50 models in principal distortion generation (not shown).
@@
Figure SI.9: Principal distortions for standard ImageNet trained and adversarially trained AlexNet and ResNet50 architectures for example base images.
@@
Figure SI.10: Example principal distortions for ResNet50 vs. AlexNet comparison where the pixelwise min/max value for the perturbation has not been constrained to avoid clipping once the perturbation is scaled. Similar to the results in the main text, we see that the principal distortion to which AlexNet is more sensitive is focused on the part of the image with high-contrast texture features, while ResNet50 is more sensitive to the perturbation targeting relatively smooth parts of the image.